OpenStudy (anonymous):

Simplify the given expression. square root negative 10 open parentheses 11 plus square root negative 5

OpenStudy (anonymous):

\[\sqrt{-10} \left( 11+\sqrt{-5} \right)\]

OpenStudy (anonymous):